### v2

# A robust and versatile deep learning model for prediction of the arterial input function in dynamic small animal FDG PET imaging

Abstract
Dynamic positron emission tomography (PET) and kinetic modeling are pivotal in advancing tracer development research in small animal studies. Accurate kinetic modeling requires precise input function estimation, traditionally achieved through arterial blood sampling. However, arterial cannulation in small animals, such as mice, involves intricate, time-consuming, and terminal procedures, precluding longitudinal studies. This work proposes a non-invasive, fully convolutional deep learning-based approach (FC-DLIF) to predict input functions directly from PET imaging data, which may eliminate the need for arterial blood sampling in the context of dynamic small-animal PET imaging.
The proposed FC-DLIF model consists of a spatial feature extractor that acts on the volumetric time frames of the dynamic PET imaging sequence, extracting spatial features. These are subsequently further processed in a temporal feature extractor that predicts the arterial input function. The proposed approach is trained and evaluated using images and arterial blood curves from data using cross validation. Further, the model applicability is evaluated on imaging data and arterial blood curves collected using two additional radiotracers (, and ). The model was further evaluated on data truncated and shifted in time, to simulate shorter, and shifted, PET scans.
The proposed FC-DLIF model reliably predicts the arterial input function with respect to mean squared error and correlation. Furthermore, the FC-DLIF model is able to predict the arterial input function even from truncated and shifted samples. The model fails to predict the AIF from samples collected using different radiotracers, as these are not represented in the training data.
Our deep learning-based input function offers a non-invasive and reliable alternative to arterial blood sampling, proving robust and flexible to temporal shifts and different scan durations.
1 Introduction
Dynamic positron emission tomography (PET) plays a critical role in the imaging of small animals, enabling in vivo visualization of tracer uptake over time, and subsequent quantification of tracer transport or binding through kinetic modeling [1]. This is actively used as a tool for downstream tasks, for example in the development of new tracers, drugs, diagnostic procedures, and disease therapies [2, 3, 4]. However, a prerequisite for kinetic modeling is knowledge of the arterial tracer concentration over time, commonly referred to as the arterial input function (AIF) [5].
Arterial blood sampling is considered the gold standard for AIF acquisition, but in preclinical settings it presents serious limitations. The procedure is technically complex, time-consuming, and terminal in mice due to the invasiveness of carotid cannulation. Furthermore, only a limited amount of blood volume can be withdrawn without altering animal physiology [6, 7]. These limitations make high-throughput and longitudinal studies infeasible and may further raise ethical concerns regarding excessive animal usage [8].
This has driven significant interest in non-invasive AIF estimation strategies. Population based input functions (PBIFs) averages time-activity curves (TACs) from demographically similar subjects [9], but fail to capture individual variability and still requires at least one blood sample to scale the population curve. Image-derived input functions (IDIFs) extract TACs from vascular regions such as the left ventricle, but are prone to inaccuracies stemming from partial-volume effects, motion artifacts, and poor signal-to-noise ratios [10, 6, 11, 12, 13]. Simultaneous estimation (SIME) approaches reduce the reliance on blood sampling by fitting both tissue kinetics and the AIF from multiple regions concurrently [14], but they require a predefined functional form for the AIF and still depend on at least one late-time blood sample to anchor the estimated curve to absolute tracer concentrations [15, 16, 17, 18].
To address these limitations, several data-driven alternatives have been proposed. In our previous research, we developed machine learning-based methods for AIF estimation using extracted TACs as input [19, 20]. While these methods eliminated the need for blood sampling, they relied on manually delineated regions of interest (ROIs) and post-imaging TAC extraction, limiting their scalability and integration into automated pipelines.
In more recent studies, it has been demonstrated that deep learning-based methods outperform more traditional machine learning-based methods bypassing the need for handcrafted features and manual ROI delineation by prediction of the AIF directly from the dynamic PET image volume [21, 22, 23, 24]. However, most of these methods adopt hybrid architectures that combine 3D convolutional layers with fully connected or recurrent layers. These architectural choices introduce key limitations. For example, fully connected layers necessitate fixed-length inputs, forcing all dynamic scans to be padded, truncated, or interpolated to a uniform number of time frames prior to inference [21, 22]. This increases computational overhead and may introduce interpolation artifacts, especially when acquisition protocols differ in temporal resolution or duration.
Recurrent-style models, such as LSTM-based networks [21, 24], attempt to model temporal dependencies explicitly, but often learn features tied to specific time indices, reducing robustness to timing shifts or variation in tracer arrival across subjects. Furthermore, several methods include post hoc fitting or model parameter regression [24, 23], adding assumptions and complexity to the pipeline. These constraints can hinder generalization and reduce flexibility in practical settings.
In this study, we propose a novel approach—fully convolutional deep learning-based input function prediction (FC-DLIF)—that overcomes these limitations by using only convolutional operations over both spatial and temporal dimensions. The fully convolutional design allows FC-DLIF to process input sequences of arbitrary length, eliminating the need for dense or recurrent layers, and removing dependence on fixed input shapes. FC-DLIF takes reconstructed 4D dynamic PET data (t, x, y, z) and directly predicts an AIF output, with no need for manual ROI segmentation, TAC extraction, temporal resampling, or post hoc fitting. By treating time as a learnable axis and detecting temporal patterns such as peak onset, plateau, and washout tails, regardless of absolute frame positions, FC-DLIF is inherently robust to timing variability and flexible across different imaging protocols. The architecture provides a streamlined, end-to-end pipeline for non-invasive AIF estimation suitable for diverse preclinical study designs.
We evaluate FC-DLIF on a dynamic PET dataset of mice with paired arterial blood data, and show that our method generalizes across protocols with time-shifted inputs and different scan durations. By accurate, non-terminal, and non-invasive AIF estimation, FC-DLIF has the potential to reduce animal use, facilitating longitudinal studies, and simplify the workflow for kinetic-modeling, aligning with the 3Rs principles of animal research [8].
2 Materials and methods
This section describes the data, the design of the proposed deep learning model, and its training procedure. Further details on data collection are provided in Appendix A.
2.1 Dataset
[…]
2.2 Model architecture
[…]
2.3 Training procedure
The ADAM optimizer [28, 29] with standard settings was used to minimize the weighted mean squared error (wMSE) between the predicted DLIF and the ground truth. The error on each time step was weighted to account for the imbalance between three parts of the curve: the first (peak), middle (intermediate), and last (tail) frames were associated with weights respectively. Training was performed with a learning rate of for epochs. -fold cross validation allowed the evaluation of the model performance over the whole training dataset. For each fold, runs were repeated for statistical rigor. Data augmentation with additive random Poisson noise injection were used during training, with the goal of exposing the model to different signal-to-noise ratios and accustom it to lower image qualities. The degraded image was generated by sampling a scalar , then, for a dynamic PET image with voxel intensities , the corresponding image with noise, is defined as:
2.4 Evaluation metrics
The predicted DLIF curves were compared time frame by time frame with the respective measured AIF using a paired t-test (), and orthogonal regression to account for measurement errors in both predicted and measured variables. Normality was assessed using quantile-quantile plots, reporting the Pearson correlation to describe the spread of the predicted curves. In order to assess the efficacy of the proposed method for downstream tasks, such as kinetic modeling, the graphical Patlak plot [30, 31] is used to produce net influx rates Ki given the reference and predicted input functions for each voxel over time. Prior to kinetic modeling, both the AIF and DLIF curves were converted to plasma input functions following [32]. DLIF-based influx estimates were further compared to those obtained from the measured AIF using the same evaluation methods as for the predicted curves.
Comparisons were done against the DLIF model proposed by Kuttner et al. [22], which will be referred to as the baseline in the following sections. To ensure fair comparisons, the baseline was trained on the same data as the FC-DLIF, also using a -fold cross validation setting.
Additionally and unrelated to the above comparisons, the FC-DLIF network’s ability to operate on data with varying temporal dimensions were explored using two tests. First, a time-shifted sample was simulated by prepending the initial time frame to the dynamic PET image—before the radiotracer uptake phase had begun—to reveal any learned time-specific dependencies. For instance, the network may insist that time frame always contains the peak of the input function during inference.
[…]
To investigate the latent representations extracted from the SFE, the t-SNE [33] algorithm was used. The algorithm was applied to one model from a single run of a specific fold (fold 2, run 9) to visualize common traits of the condensed spatial representation.
3 Results
First, this section summarizes the results of the proposed FC-DLIF model during cross-validation on our dataset when compared against the baseline, both in terms of errors on AIF estimation and, consequently, after tracer kinetic modeling. Then, some examples of AIF estimation show the performance and versatility of the proposed model with respect to the input data. Finally, the features extracted by the model’s SFE is visualized and inspected using the t-SNE algorithm [33].
3.1 Quantitative results
Fig. 2 summarizes the distribution of mean squared error (MSE, Fig. 2(a)) and injected noise (Fig. 2(b)) across all samples. FC-DLIF consistently achieved lower MSE values than the baseline, with a noticeably narrower spread and a more favorable distribution across percentiles. In terms of noise, FC-DLIF also demonstrated reduced variability, being less sensitive to noise in the input data.
In particular, the proposed model has close to of MSE vales lower than the median of the baseline model, and generally produce predictions more aligned with the AIF, as seen in the percentile (best of predictions), which is slightly lower than the baseline (Fig. 2(a)). These results are further consolidated when comparing the same metrics over temporal regions, as shown in Fig. 9, where the proposed method exhibits less bias in most segments.
In Fig. 3, the predicted AIF values from each model are compared against ground truth measurements on a point-wise basis. The proposed method achieves higher and Pearson correlation values, with predictions more closely aligned to the ideal line, indicating a tighter spread of predictions around the regression line. While the baseline exhibits systematic deviations, particularly overestimating intermediate and tail values and underestimating peak amplitudes, FC-DLIF maintains a tighter spread around the regression line.
Predictions from the baseline model are generally overestimating small values (less than ), and underestimating larger values, associated with the peak of the input function curve. The same characteristics are not seen for the FC-DLIF method, but instead the spread is generally larger above values of , which is also shown in Fig. 9, for the peak labeled segments.
Downstream implications of AIF estimation accuracy are reflected in kinetic modeling results shown in Fig 4, where voxel-wise values derived from each model are compared to reference values obtained using the measured AIF and the Patlak model [34, 35]. FC-DLIF again yields higher agreement with the reference, confirming that improvements in input function estimation translate to better physiological parameter estimation.
Results on the unseen tracer data ( and ) showed decreased performance for both models, as expected due to tracer-specific uptake characteristics. Additional scatterplots and analysis are provided in the Appendix B (Figs. 10 and 11).
3.2 Qualitative results
This section evaluates the predictions of the proposed FC-DLIF model. This is followed by an experiment that shift the PET images in time, to delay tracer injection start. Then the PET images are truncated, which simulates simultaneous tracer injection and scan start.
Fig. 5 shows three examples of input functions predicted by the proposed FC-DLIF model. The mean curve and standard deviation of the model’s predictions are included for the best (Fig. 5(a)), median (Fig. 5(b)), and worst (Fig. 5(c)) sample according to the MSE.
Fig. 6 displays the versatility of the proposed model owing to its fully convolutional design. In particular, the example on the left is the output of the model when the input data is time-shifted by prepending the first frame of the dynamic PET image to itself. The example on the right is the result obtained by truncating the input both at the beginning and at the end of the time series, removing several time steps.
Finally, t-SNE is used to visualize the similarities between the high-dimensional vectors extracted by the SFE for each time step of each mouse, on both the training and test datasets. Fig. 7 and 8 illustrate the results of the t-SNE dimensionality reduction for the groups of injected volume and injected tracers respectively. To the left, the colormap indicates the peak, mid, and tail time steps in shades of red, green, and blue respectively. On the right hand side, the colormap indicate certain attributes—the amount of injected in Fig. 7, and what tracer is used in Fig. 8.
Fig. 7 shows a clear temporal progression in feature space. Early frames with low uptake cluster tightly on the left, followed by a gradual arc through peak and post-distribution frames. The pattern reflects the biological progression of tracer dynamics and suggests that the SFE captures consistent spatial characteristics across time.
Likewise, Fig. 8 shows the t-SNE projection, from different tracer groups, that were kept aside during training, comprised of , , and injected mice. While the early and peak phases of all tracers appear to follow a similar initial trajectory, clear deviations form in the later frames. In particular, the steady-state representations of the and groups form separate clusters, distinct from the tail-phase points. This indicates that the model’s spatial encoder captures consistent tracer-specific signatures not seen during training. Additionally, the group also diverges around the late peak phase, forming a distinct cluster that may explain the model’s reduced accuracy for this group. These observations align with the systematic under- and overestimation patterns seen in the early and late phases of predicted AIFs for these out-of-distribution tracers.
4 Discussion
This study introduces FC-DLIF, a robust and flexible deep learning model for the non-invasive estimation of AIF directly from dynamic PET imaging data. Compared to existing methods, FC-DLIF demonstrates superior accuracy, reduced bias, and improved robustness to noise and input variability, while maintaining flexibility with respect to temporal shifts and varying scan lengths. The fully convolutional design enables the model to generalize across different imaging protocols without the need for rigid preprocessing or fixed input dimensions.
4.1 Quantitative analysis
Fully convolutional design leads to consistent improvements
Across all quantitative evaluations, FC-DLIF demonstrated superior accuracy compared to the baseline model. While both methods use 3D convolutional backbones to extract spatial features, the fully convolutional design of FC-DLIF, that comprises separate spatial and temporal feature extractors, appears to contribute to more accurate AIF predictions. In particular, each time frame is processed independently through the same spatial filters in the SFE, potentially reducing time-point-specific biases during feature extraction. The subsequent TFE, built with 1D convolutions (Fig. 1), is then able to learn temporal dependencies over a time-series of uniformly extracted spatial representations.
Although the precise source of performance gain is difficult to isolate, the two-step network inherently enforces a type of implicit regularization. This separation of concerns allows each subnetwork to specialize; spatial encoding in the SFE; and temporal modeling in the TFE, potentially introducing a form of structural regularization. This modular architecture may help the network learn more generalized temporal patterns in tracer dynamics, leading to improved alignment with the true AIF. The benefits are clearly reflected in the lower MSE distribution (Fig. 2(a)), being less affected by a noisy input signal (Fig. 2(b)), and a tighter correspondence between predicted and true AIF values (Fig. 3).
Improved downstream kinetic modeling
The improvements in AIF prediction also translate into better physiological parameter estimation, as evidenced by the voxel-wise comparisons (Fig. 4), and region-wise comparisons (Fig. 13(e) and Fig. 14(e)). Accurate AIF estimation is critical for computing reliable kinetic parameters using models such as Patlak [34, 35] and compartment models [36], and FC-DLIF’s enhanced performance further supports its potential to replace invasive blood sampling without compromising modeling accuracy.
Generalization challenges with unseen tracers
When applied to a dataset using tracers not seen during training ( and tracers), both models exhibited reduced performance. This result is consistent with the expectation that different tracers produce distinct uptake patterns and temporal kinetics, making generalization difficult without exposure during training. Future work may explore strategies for improving generalization across tracers, such as domain adaptation [37, 38], tracer-aware conditioning, or semi-supervised strategies [39].
4.2 Qualitative analysis
Understanding prediction characteristics
To further gain insight into the model’s behavior, the AIF predictions from FC-DLIF were examined over a range of samples (Fig. 5). Even in the worst-performing example, the model captures the general shape of the AIF, particularly during the post-distribution phase. This aligns with the scatter-based results from Fig. 3, where the largest errors are observed around the peak uptake frames, while the steady-state phase shows more consistent alignment with ground truth. This discrepancy likely stems from the temporal characteristics of the input data itself: the uptake phase is captured in only a few frames, and exhibits high inter-frame variability, making it difficult to learn reliably. In contrast, the tail of the curve, corresponding to tracer equilibrium, spans more frames and features more consistent signal characteristics, yielding more stable predictions (as seen in Fig. 9).
Importantly, for downstream kinetic modeling tasks such as Patlak analysis [34, 35], the steady-state segment of the AIF is the only part used to derive macro-parameters like . This suggests that even if the peak predictions are imperfect, FC-DLIF’s accurate modeling of the linear phase still enables reliable kinetic quantification, when using the Patlak model.
Robustness to shifts and truncation
One of the practical advantages of FC-DLIF is its architectural flexibility, which is demonstrated by the model’s robustness to time shifts and partial input sequences, a trait which has not been seen in other similar works. When the input time series was artificially shifted or truncated, the model continued to produce consistent AIF predictions (Fig. 6). This behavior stems directly from the model’s fully convolutional structure, which applies the same spatial and temporal filters regardless of absolute time index or input length, emphasizing its practicality in real-world settings, where scan durations and start times may vary between protocols or subjects.
What the model learns from spatial features
To better understand the internal representations learned by FC-DLIF, the spatial features extracted by the SFE were visualized using t-SNE [33]. Across all mice and time points, the t-SNE component projection reveals a coherent temporal progression: from early frames with low uptake, through peak uptake, to the post-distribution tail. This suggests that the SFE successfully encodes temporally meaningful spatial information, even though it processes each frame independently. Notably, these feature clusters were consistent across both training and test sets, with no clear separation, indicating good generalization and successfully not overfitting. Furthermore, the internal feature space showed no meaningful correlation with injected volume, suggesting that the SFE learned to abstract away from trivial scaling differences in tracer volume, which is an encouraging sign of biological relevance and robustness (Fig. 7).
Out-of-distribution tracers reveal biological misalignment
Despite showing reasonable results and potential for improvements, t-SNE projections of mice injected with unseen tracers ( and ) revealed key deviations. While their early and peak-phase representations partially aligned with the training set, their tail-phase features diverged and formed distinct clusters. This reflects fundamental differences in tracer kinetics and binding properties [40]. Correspondingly, FC-DLIF struggled to predict accurate AIFs for these tracers, typically underestimating peak intensity and overestimating post-distribution activity. This mismatch to some extent mirrors the temporal drift observed in the feature space, and suggests that the model’s learned temporal filters in the TFE are not readily transferable across tracer types without retraining.
Spatial dimensions limitations
While the proposed method accepts input data of varying temporal dimensions, a pretrained model constrains future studies to preprocess their spatial dimensions to match that of the data it was trained on. Since transformations and interpolations over spatial dimensions are common and easier to visualize than temporal ones, this is not seen as a large hurdle.
Toward generalization through transfer learning
Despite these limitations, the structure of the learned feature space provides an opportunity for targeted adaptation. Since the SFE appears to produce meaningful spatial embeddings even for unseen tracers, it may be possible to freeze this component and retrain only the temporal extractor (TFE) on a small amount of new tracer data, to map the distinct clusters back onto the common feature manifold. This strategy aligns with principles from transfer learning and domain adaptation [37, 38], and could substantially reduce the data requirements for generalizing FC-DLIF to new tracers. Given that most variation across tracers lies in their temporal behavior, this modular approach, reusing spatial filters while tailoring the temporal decoder, offers a promising path forward.
Compact design for efficient deployment
A final yet key advantage of FC-DLIF is its compact size. The model end-to-end processes full 4D PET volumes yet comprises only trainable parameters (352 KB). This efficiency arises from its internal structure, where each 3D time frame is processed independently using shared spatial filters in the SFE, followed by a lightweight 1D convolutional TFE consisting of 4 layers. For comparison, the input images are approximately 71 MB (64-bit floating point) each, making the model over 200 times smaller than a single input scan. Despite its modest footprint, which is comparable to the parameter (360 KB) baseline model, FC-DLIF achieves superior predictive performance, making it an attractive candidate for real-time or embedded applications.
5 Conclusion
This study introduces a non-invasive, deep learning-based input function predictor, that provides a flexible method that is invariant to time-shifting and truncation of the sequence length. Altogether, FC-DLIF represents a promising tool for preclinical PET studies, removing the need for invasive blood sampling and enabling more efficient and scalable experimental designs.
Acknowledgments
This work is funded through Visual Intelligence, a Center for Research-based Innovation funded by the Research Council of Norway, grant no. 309439 and 303514. It was further supported by Tromsø Research Foundation (19_PET-NUKL) through the 180∘N Norwegian Nuclear Medicine Consortium and UiT Innovation funds (UiT Talent).
Appendix A Data acquisition and processing
A.1 Animal experiments
The imaging data of mice used to train and evaluate the FC-DLIF model were acquired at UiT The Arctic University of Norway (UiT). All animal experiments were approved by the Norwegian Food Safety Authority; FOTS id 29689. In total, healthy female mice from three different strains were included in this study: BALB/cJRj (N = ), C57BL/6JRj (N = ) and Balb/cAnNCrl (N = ), where BALB/cJRj made up the main group of samples. The remaining were set aside for evaluating the effects of the tracers (N = ) and (N = ). The mice were weeks of age upon arrival to the UiT animal facility, and fed ad libitum a standard rodent diet.
A.2 Arterial and venous cannulation during image acquisition
At the time of PET/CT imaging, the mouse age was , with a corresponding weight of . The study used three different radiotracers: , , and . The mice receiving ( animals) were fasted for prior to injection, whereas the remaining animals—given and —were not fasted. The mice were anesthetized prior to the scan, weighed, and placed on a heated plate () while receiving oxygen through a mask. A venous catheter was inserted into the tail vein of each mouse for radiotracer injection. The blood glucose during venous cannulation was measured to . An incision in the neck enabled surgical cannulation of the carotid artery, allowing blood to be routed through a radiation detector at a withdrawal rate of . This setup facilitated concurrent measurements of whole blood activity during the PET scan with a temporal resolution of . To enable continuous arterial line measurements, an arterial–venous shunt was established, forming a closed loop between arterial sampling and venous reinjection. Blood flowed sequentially from the arterial sampling line, radiation detector, a peristaltic pump, and a Y-connector—enabling intravenous radiotracer injection—before being reinfused via the venous injector, thereby preventing excessive blood loss.
[…]
A.3 Calibration and AIF processing
To allow for delay correction of the manual blood sample measurements in the arteriovenous shunt, the time delay between the radiation detector and the Y-connector was measured during the first pass of arterial blood to . The continuous line radiation detector was calibrated with three manual blood samples taken in a late stage or post-scan by measuring blood dripped from the arterial catheter over . This also enabled measurement of the actual arterial withdrawal rate during each mouse scan.
A calibration factor for the continuous arterial line measurements was derived as the ratio of the average signal from the continuous line measurements and the manual blood samples collected during the same blood sampling at each time point, corrected for delay. Calibration factors outside three scaled median absolute deviations from the median factors were considered outliers and discarded [41]. An overall calibration factor was determined as the average factor from the included blood sample factors. The arterial input function (AIF) for each mouse was obtained by scaling the continuous line measurement data by this average calibration factor.
A.4 Image reconstruction and processing
[…]
Appendix B Additional results
B.1 Detailed results for individual imaging phases
Figure 9 visualizes the distribution of error between the predicted input function from the baseline model [22] and FC-DLIF compared with the AIF, as a function of time since the start of tracer injection. A predominantly negative distribution in the plot would indicate the model underestimating the AIF.
B.2 Results on unseen tracers dataset
Figure 10 compares the baseline [22] and the proposed FC-DLIF models’ SUV predictions with the measured AIF. The lines are computed as the orthogonal regression line fitted to the point distributions of each respective color. A perfect predictor would have SUV points distributed across the black, dashed, line.
Figure 11 uses the predicted input function from either model in a Patlak graphical analysis [34, 35] to derive the kinetic parameters .
Figure 12 shows examples of input functions predicted by the proposed FC-DLIF model. The mean curve and the standard deviation over the runs are depicted for the best (left), median (middle), and worst (right) sample according to the mean squared error (MSE).
B.3 Detailed kinetic modeling results
Further kinetic modeling results using the predicted input functions from both models are summarized in Figures 13 and 14. The kinetic parameters were estimated using an irreversible two-tissue compartment model [36] in two different regions: myocardium and brain. The estimated parameters were compared with those obtained using the measured AIF and DLIF. The quantile–quantile plots show the distribution of the kinetic parameters estimated using orthogonal regression.
Compared to the Patlak analysis presented in the main manuscript, the kinetic parameters derived from the two-tissue compartment model are less stable because the full measurement period is used in the fitting, and not only the linear phase. Outliers are therefore removed if they lie more than three standard deviations away from the mean of each parameter distribution. Once an outlier is detected, it is removed from all parameters in the respective region for consistency. This results in and samples remaining for the myocardium and brain regions, respectively.
The observed myocardium kinetic parameters (Figures 13 and 13(e)) follow the reference distribution well, with larger spread in , , and , as well as outliers not removed by the three standard deviation rule in and the blood volume fraction. Influx rates (Figure 13(e)) were higher than previously reported [42, 43]. This is likely due to shorter fasting times in our study, where longer fasting times are known to reduce influx rates in the myocardium [42, 43].
The parameter distributions in the brain (Figure 14) had fewer outliers, and similar to the myocardium, , , and had larger spread than the uptake and influx rate. The ranges of the brain kinetic parameters are in line with previous literature [44].